# ICLR: In-Context Learning of Representations

### v1

@@ 7 Discussion
Limitations. We do emphasize that our work has a few limitations. Namely, PCA, or more broadly, low dimensional visualizations of high dimensional data can be difficult to interpret or sometimes even misleading. Despite such difficulties, we provide theoretical connections between energy minimization and principal components to provide a compelling explanation for why structures elicited via PCA faithfully represent the in-context graph structure. Second, we find a strong, but nevertheless incomplete, causal relationship between the representations found by PCA and the model’s predictions. We view the exact understanding of how these representations form, and the exact relationship between the representations and model predictions as an interesting future direction, especially given that such underlying mechanism seems to depend on the scale of the context.
@@ Acknowledgments
We greatly thank the National Deep Inference Fabric (NDIF) pilot program (Fiotto-Kaufman et al., 2024), especially, Emma Bortz, Jaden Fiotto-Kaufman, Adam Belfki, David Bau, and the team who provided us with access to representations of Llama models, making this work possible. CFP and HT acknowledge the support of Aravinthan D.T. Samuel, Cecilia Garraffo and Douglas P. Finkbeiner. CFP, ESL, KN, MO, and HT are supported by the CBS-NTT Program in Physics of Intelligence. MW acknowledges support from a Superalignment Fast Grant from OpenAI, Effective Ventures Foundation, Effektiv Spenden Schweiz, and the Open Philanthropy Project. Part of the computations in this paper were run on the FASRC cluster supported by the FAS Division of Science Research Computing Group at Harvard University. ESL thanks Eric Bigelow for crucial discussions that helped define several hypotheses pursued in this work, and the Harvard CoCoDev lab (especially Peng Qian) and Talia Konkle for feedback on an earlier version of the paper. CFP thanks Zechen Zhang, Eric Todd, Clement Dumas, and Shivam Raval for useful discussions. All authors thank David Bau and his lab for useful feedback on the paper’s results.
@@ Appendix B The Connection Between Energy Minimization and PCA Stucture
then the $k$-th principle component of ${\bm{H}}$ (for $k\in[r]$) is be ${\bm{z}}_{k+1}$.
@@
Since ${\bm{u}}_{k}$’s are singular vectors, we have ${\bm{u}}_{k}$’s are orthogonal to each other. Using Theorem 1 in Fan (1949), we know that for any $s^{\prime}\in[n]$, the minimizer of $\sum_{k=1}^{s^{\prime}}E_{\mathcal{G}}({\bm{u}}_{k})$ is ${\bm{u}}_{k}={\bm{z}}_{k},\forall k\in[s^{\prime}]$. Therefore, it is evident that the minimizer of $\sum_{k=1}^{s}\sigma_{k}^{2}E_{\mathcal{G}}({\bm{u}}_{k})$ must satisfies ${\bm{u}}_{k}={\bm{z}}_{k},\forall k\in[s]$, since from the above argument of $\sigma_{k}$’s and the given condition condition we know that $\sigma_{1}>\sigma_{2}>\cdots>\sigma_{s}>0$.
Now we have proved that ${\bm{u}}_{k}={\bm{z}}_{k},\forall k\in[s]$. Next we consider the output of PCA. Let ${\bm{p}}_{k}$ be the $k$-th principle component output by the PCA of ${\bm{H}}$. We know that ${\bm{p}}_{k}$ is the eigenvector of
@@
which, again from Spectral Theorem, is the eigenvector of the second largest eigenvalue of ${\bm{H}}{\bm{H}}^{\top}$, which is ${\bm{u}}_{2}={\bm{z}}_{2}$. Using an induction and the same reasoning, it follows that for any $k\in[s]$, we have ${\bm{p}}_{k}={\bm{z}}_{k+1}$. This proves the proposition. ∎

### v2

@@ Acknowledgments
We greatly thank the National Deep Inference Fabric (NDIF) pilot program (Fiotto-Kaufman et al., 2024), especially, Emma Bortz, Jaden Fiotto-Kaufman, Adam Belfki, David Bau, and the team who provided us with access to representations of Llama models, making this work possible. CFP and HT acknowledge the support of Aravinthan D.T. Samuel, Cecilia Garraffo and Douglas P. Finkbeiner. CFP, ESL, KN, MO, and HT are supported by the CBS-NTT Program in Physics of Intelligence. AL and MW acknowledges support from the Superalignment Fast Grant from OpenAI. MW also acknowledges support from Effective Ventures Foundation, Effektiv Spenden Schweiz, and the Open Philanthropy Project. Part of the computations in this paper were run on the FASRC cluster supported by the FAS Division of Science Research Computing Group at Harvard University. ESL thanks Eric Bigelow for crucial discussions that helped define several hypotheses pursued in this work, and the Harvard CoCoDev lab (especially Peng Qian) and Talia Konkle for feedback on an earlier version of the paper. CFP thanks Zechen Zhang, Eric Todd, Clement Dumas, and Shivam Raval for useful discussions. All authors thank David Bau and his lab for useful feedback on the paper’s results.
@@ Appendix B The Connection Between Energy Minimization and PCA Stucture
then the $k$-th principle component of ${\bm{H}}$ (for $k\in[r]$) is ${\bm{z}}^{(k+1)}$.
@@
Since ${\bm{u}}_{k}$’s are singular vectors, we have ${\bm{u}}_{k}$’s are orthogonal to each other. Using Theorem 1 in Fan (1949), we know that for any $s^{\prime}\in[n]$, the minimizer of $\sum_{k=1}^{s^{\prime}}E_{\mathcal{G}}({\bm{u}}_{k})$ is ${\bm{u}}_{k}={\bm{z}}^{(k)},\forall k\in[s^{\prime}]$. Therefore, it is evident that the minimizer of $\sum_{k=1}^{s}\sigma_{k}^{2}E_{\mathcal{G}}({\bm{u}}_{k})$ must satisfies ${\bm{u}}_{k}={\bm{z}}^{(k)},\forall k\in[s]$, since from the above argument of $\sigma_{k}$’s and the given condition condition we know that $\sigma_{1}>\sigma_{2}>\cdots>\sigma_{s}>0$.
Now we have proved that ${\bm{u}}_{k}={\bm{z}}^{(k)},\forall k\in[s]$. Next we consider the output of PCA. Let ${\bm{p}}_{k}$ be the $k$-th principle component output by the PCA of ${\bm{H}}$. We know that ${\bm{p}}_{k}$ is the eigenvector of
@@
which, again from Spectral Theorem, is the eigenvector of the second largest eigenvalue of ${\bm{H}}{\bm{H}}^{\top}$, which is ${\bm{u}}_{2}={\bm{z}}^{(2)}$. Using an induction and the same reasoning, it follows that for any $k\in[s]$, we have ${\bm{p}}_{k}={\bm{z}}^{(k+1)}$. This proves the proposition. ∎
@@ C.7 Additional Empirical Verifications of Transition Predictions
Appendix D Limitations
We emphasize that our work has a few limitations. Namely, PCA, or more broadly, low dimensional visualizations of high dimensional data can be difficult to interpret or sometimes even misleading. Despite such difficulties, we provide theoretical connections between energy minimization and principal components to provide a compelling explanation for why structures elicited via PCA faithfully represent the in-context graph structure. Second, we find a strong, but nevertheless incomplete, causal relationship between the representations found by PCA and the model’s predictions. We view the exact understanding of how these representations form, and the exact relationship between the representations and model predictions as an interesting future direction, especially given that such underlying mechanism seems to depend on the scale of the context.